### v1

# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

Abstract
Many applications, e.g., in shared mobility, require coordinating a large number of agents. Mean-field reinforcement learning addresses the resulting scalability challenge by optimizing the policy of a representative agent. In this paper, we address an important generalization where there exist global constraints on the distribution of agents (e.g., requiring capacity constraints or minimum coverage requirements to be met). We propose Safe-$\text{M}^{3}$-UCRL, the first model-based algorithm that attains safe policies even in the case of unknown transition dynamics. As a key ingredient, it uses epistemic uncertainty in the transition model within a log-barrier approach to ensure pessimistic constraints satisfaction with high probability. We showcase Safe-$\text{M}^{3}$-UCRL on the vehicle repositioning problem faced by many shared mobility operators and evaluate its performance through simulations built on Shenzhen taxi trajectory data. Our algorithm effectively meets the demand in critical areas while ensuring service accessibility in regions with low demand.
1 Introduction
Multi-agent reinforcement learning (MARL) is a rapidly growing field that seeks to understand and optimize the behavior of multiple agents interacting in a shared environment. MARL has a wide range of potential applications, including vehicle repositioning in shared mobility services (e.g., moving idle vehicles from low-demand to high-demand areas [35]), swarm robotics (e.g., operating a swarm of drones), and smart grids (e.g., operating a network of sensors in electric system [46]). The interactions between agents in these complex systems introduce several challenges, including non-stationarity, scalability, competing learning goals, and varying information structure.
Mean-Field Control (MFC) addresses the scalability and non-stationarity hurdles associated with MARL by exploiting the insight that many relevant MARL problems involve a large number of very similar agents. Instead of focusing on the individual agents and their interactions, MFC considers an asymptotically large population of identical agents and models them as a distribution on the state space. The solutions obtained by MFC are often sufficient for the finite-agent equivalent problem [29, 40] in spite of the introduced approximations. An example of such a system is a ride-hailing platform that serves on-demand trips with a fleet of vehicles. The platform needs to proactively reposition idle vehicles based on their current locations, the locations of the other vehicles in the fleet, and future demand patterns (see Figure 1) to maximize the number of fulfilled trips and minimize customer waiting times. Additionally, the platform may be obligated by external regulators to guarantee service accessibility across the entire service region. The problem quickly becomes intractable as the number of vehicles increases.
In this paper, we focus on learning the safe optimal policies in a setting where the underlying dynamics are unknown. In most real-world systems, the dynamics must be learned from the data obtained from repeated interactions with the environment. We assume that the cost of obtaining data from the environment is high and seek to design a model-based solution that efficiently uses the collected data. Existing works consider solving the MFC problem via model-free or
Figure 1: An illustrative example of taxis’ spatial distribution (light-blue scatters), repositioning trips (blue arrows), and a taxi’s trajectory as a sequence of passenger trips (red arrows).
model-based methods without safety guarantees. However, the proposed Safe Model-Based Multi-Agent Mean-Field Upper-Confidence RL (Safe-$\text{M}^{3}$-UCRL) algorithm focuses on learning optimal policies for MFC while avoiding undesired or dangerous distributions of agents’ population.
Contributions. Section 3 considers a novel mean-field setting with safety constraints and defines a novel comprehensive notion of global population-based safety. To address safety-constrained environments featuring a large population of agents, in Section 4, we propose a model-based mean-field reinforcement learning algorithm called Safe-$\text{M}^{3}$-UCRL. Our algorithm leverages epistemic uncertainty in the transition model, employing a log-barrier approach to guarantee pessimistic satisfaction of safety constraints and enables the derivation of safe policies. In Section 5, we conduct empirical testing on the vehicle repositioning problem, a challenge commonly faced by shared mobility operators. Our results demonstrate that the policies derived using Safe-$\text{M}^{3}$-UCRL successfully fulfill demand in critical demand hotspots while ensuring service accessibility in areas with lower demand.
2 Related Work
Our notion of safety for the mean-field problem extends the frameworks of Mean-Field Games (MFG) and Mean-Field Control (MFC) [31, 32, 26, 25]. For a summary of the progress focusing on MFGs, see [33] and references therein. We focus on MFC in this work which assumes cooperative agents in contrast to MFGs which assume competition. [19, 41, 22, 23, 5] address the problem of solving MFCs with known dynamics, i.e., planning, while [52, 9, 4, 3, 8, 53, 10, 51] consider model-free Q-Learning and Policy Gradient methods in various settings. Closest to our approach, [43] introduces $\text{M}^{3}$-UCRL, a model-based, on-policy algorithm, which is more sample efficient than other proposed approaches. Similarly to [28, 12, 15] for model-based single-agent RL and [47] for model-based MARL, $\text{M}^{3}$-UCRL uses the epistemic uncertainty of the transition dynamics model to design optimistic policies that efficiently balance exploration and exploitation in the environment and maximize sample efficiency. This is also the setting of our interest. However, there is no consideration of safety in any of these methods.
In terms of safety, there are two main ways of handling it in RL [18]; assigning significantly lower rewards to unsafe states [38] and providing additional knowledge to the agents [49] or using the notion of controllability to explicitly avoid unsafe policies [20]. Furthermore, the following approaches combine the two methods; [6] uses Lyapunov functions to restrict the safe policy space, [11] projects unsafe policies to a safe set via a control barrier function, and [2] introduces shielding, i.e., correcting actions only if they lead to unsafe states. The aforementioned works consider only individual agents, while in large-scale multi-agent environments, maintaining individual safety becomes intractable, and the focus shifts towards global safety measures. [57] show how to use the notion of returnability to define a safe, decentralized, multi-agent version of Q-learning, however, their approach does not scale well for large systems. [40] introduces constraints to the MFC by defining a cost function and a threshold that the discounted sum of costs can not exceed. We propose a different formulation that restricts the set of feasible mean-field distributions at every step, therefore, allowing for more specific control over constraints and safe population distributions.
3 Problem Statement
We consider the Mean-Field Control (MFC) model to capture a collective behavior of a large number of collaborative agents operating in the shared stochastic environment. This model assumes the limiting regime of infinitely many agents and homogeneity. Namely, all agents are identical and indistinguishable. Solving MFC amounts to finding an optimal policy for a single, so-called, representative agent due to the identical agent assumption. The representative agent interacts with the distribution of agents instead of focusing on individual interactions and optimizes a collective reward. Its policy is used to control all the agents in the environment.
Formally, we consider the episodic setting, where episodes $n=1,\ldots,N$ each have $t=0,\ldots,T-1$ discrete steps. The state space $\mathcal{S}\subseteq\mathbb{R}^{p}$ and action space $\mathcal{A}\subseteq\mathbb{R}^{q}$ are the same for every agent. We use $s_{n,t}^{(i)}\in\mathcal{S}$ and $a_{n,t}^{(i)}\in\mathcal{A}$, to denote the state and action of agent $i\in\{1,\ldots,m\}$ in episode $n$ at step $t$. At every $n$ and $t$, the mean-field distribution $\mu_{n,t}\in\mathcal{P}(\mathcal{S})$ describes the global state with $m$ identical agents when $m\rightarrow+\infty$:
$\mu_{n,t}(s)=\lim_{m\xrightarrow{}\infty}\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(s_{n,t}^{(i)}=s),$
where $\mathbb{I}(\cdot)$ is the indicator function, and $\mathcal{P}(\mathcal{S})$ is the set of probability measures over the state space $\mathcal{S}$.
We posit that the reward $r:\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}\rightarrow\mathbb{R}$ of the representative agent is known and that it depends on the states of the other agents through the mean-field distribution. [^1] Our framework easily extends to the unknown reward case by learning the reward function similarly to the unknown dynamics and estimating its epistemic uncertainty (see [12]). Before every episode $n$, the representative agent selects a non-stationary policy profile $\bm{\pi}_{n}=(\pi_{n,0},\ldots,\pi_{n,T-1})\in\Pi$ where individual policies [^2] We use the term policy profile to indicate non-stationary policies, while we refer to their components as individual policies. are of the form $\pi_{n,t}:\mathcal{S}\times\mathcal{P}(\mathcal{S})\rightarrow\mathcal{A}$ and $\Pi$ is the set of admissible policy profiles. The policy profile is then shared with all the agents that choose their actions according to $\bm{\pi}_{n}$ during episode $n$.
We consider a general family of transition dynamics $f:\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}\rightarrow\mathcal{S}$ defined as follows. Given the current mean-field distribution $\mu_{n,t}$, the representative agent’s state $s_{n,t}$ and its action $a_{n,t}$, the next representative agent’s state $s_{n,t+1}$ is given by:
$s_{n,t+1}=f(s_{n,t},\mu_{n,t},a_{n,t})+\varepsilon_{n,t},$ (1)
where $\varepsilon_{n,t}$ is the Gaussian noise with known variance. We assume that the transition dynamics are unknown and are to be inferred from collected trajectories across episodes.
Mean-field transition dynamics. State-to-state transition map in Equation 1 naturally extends to a mean-field transition dynamics induced by a policy profile $\bm{\pi}_{n}$ and dynamics $f$ in episode $n$ (see [43, Lemma 1]):
$\mu_{n,t+1}(ds^{\prime})=\int_{\mathcal{S}}\mathbb{P}[s_{n,t+1}\in ds^{\prime}]\mu_{n,t}(ds),$ (2)
where $s_{n,t+1}$ is the next representative agent state and $\mu_{n,t}(ds)=\mathbb{P}[s_{n,t}\in ds]$ under $\bm{\pi}_{n}$ for all $t$. To simplify the notation, we use $U(\cdot)$ to denote the mean-field transition function from Equation 2, i.e., we have $\mu_{n,t+1}=U(\mu_{n,t},\pi_{n,t},f)$. We further introduce the notation $z_{n,t}\in\mathcal{Z}=\mathcal{S}\times\mathcal{P}(\mathcal{S})\times\mathcal{A}$ to denote the tuple $(s_{n,t},\mu_{n,t},a_{n,t})$.
For a given policy profile $\bm{\pi}_{n}$ and mean-field distribution $\mu$, the performance of the representative agent at step $t$ is measured via its expected future reward for the rest of the episode, i.e., $\mathbb{E}\left[\sum_{j=t}^{T-1}r(s_{n,j},\mu_{n,j},a_{n,j})\rvert\mu_{n,t}=\mu\right]$. Here, the expectation is over the randomness in the state transitions as well as over the sampling of the representative agent’s initial state, i.e., $s_{n,t}\sim\mu$.
3.1 Safe Mean-Field Reinforcement Learning
We define a safety function $h:\mathcal{P}(\mathcal{S})\to\mathbb{R}$ applied to the mean-field distribution as follows. Given some hard safety threshold $C\in\mathbb{R}$, we consider mean-field distributions $\mu$ as safe if they satisfy $h(\mu)\geq C$, or, equivalently, if
$h_{C}(\mu):=h(\mu)-C\geq 0.$ (3)
We denote as $\zeta=\{\mu\in\mathcal{P}(\mathcal{S}):h_{C}(\mu)\geq 0\}$, the safe mean-field distribution set for a safety constraint $h_{C}(\cdot)$. Hence, our focus is on the safety of the system as a whole rather than the safety of individual agents, as it becomes intractable handling individual agents’ states and interactions in the case of a large population. For the sake of exposition, we use a single constraint, however, our approach is directly extendable to multiple constraints.
For a given initial distribution $\mu_{0}$, we formally define the Safe Multi-agent Mean-field Reinforcement Learning problem as follows:
$\bm{\pi}^{*}$ $=\operatorname*{arg\,max}_{\bm{\pi}\in\Pi}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(s_{t},\mu_{t},a_{t})\Big{\rvert}\mu_{0}\Bigg{]}$ (4a)
[…]
$s_{t+1}$ $=f(s_{t},\mu_{t},a_{t})+\varepsilon_{t}$ (4c)
$\mu_{t+1}$ $=U(\mu_{t},\pi_{t},f)$ (4d)
$h_{C}(\mu_{t+1})$ $\geq 0,$ (4e)
where we explicitly require induced mean-field distributions $\{\mu_{t}\}_{t=1}^{T}$ to reside in the safe set $\zeta$ by restricting the set of admissible policy profiles $\Pi$ to policy profiles that induce safe distributions. To ensure complete safety, we note that the initial mean-field distribution $\mu_{0}$ must be in the safe set $\zeta$ as our learning protocol does not induce $\mu_{0}$ (Algorithm 1) and $s_{0}$ is sampled from $\mu_{0}$.
We make the following assumptions about the environment using the Wasserstein $1$-distance defined as $W_{1}(\mu,\mu^{\prime}):=\inf_{\gamma\in\Gamma(\mu,\mu^{\prime})}\mathbb{E}_{(x,y)\sim\gamma}\|x-y\|_{1}$ where $\Gamma(\mu,\mu^{\prime})$ is the set of all couplings of $\mu$ and $\mu^{\prime}$ (i.e., a joint probability distributions with marginals $\mu$ and $\mu^{\prime}$). We further define the distance between $z=(s,\mu,a)$ and $z^{\prime}=(s^{\prime},\mu^{\prime},a^{\prime})$ as $d(z,z^{\prime}):=\|s-s^{\prime}\|_{2}+\|a-a^{\prime}\|_{2}+W_{1}(\mu,\mu^{\prime})$.
Assumption 1 (Transition dynamics Lipschitz continuity).
The dynamics function $f(\cdot)$ is $L_{f}$-Lipschitz-continuous, i.e., $\|f(z)-f(z^{\prime})\|_{2}\leq L_{f}d(z,z^{\prime})$.
Assumption 2 (Mean-field policies Lipschitz continuity).
The individual policies present in any admissible policy profile in $\Pi$ are $L_{\pi}$-Lipschitz-continuous, i.e., $\|\pi(s,\mu)-\pi(s^{\prime},\mu^{\prime})\|_{2}\leq L_{\pi}(\|s-s^{\prime}\|_{2}+W_{1}(\mu,\mu^{\prime}))$ for all $\pi\in\bm{\pi}\in\Pi$.
Assumption 3 (Reward Lipschitz continuity).
The reward function $r(\cdot)$ is $L_{r}$-Lipschitz-continuous, i.e., $\|r(z)-r(z^{\prime})\|_{2}\leq L_{r}d(z,z^{\prime})$.
These assumptions are considered standard in model-based learning [28, 12, 47, 43] and mild, as individual policies and rewards are typically designed in such a way that they meet these smoothness requirements. For example, we use neural networks with Lipschitz-continuous activations to represent our policies. Therefore the obtained policies are also Lipschitz-continuous when the weights are bounded (see Section D.3).
3.2 Statistical Model and Safety Implications
The representative agent learns about the unknown transition dynamics by interacting with the environment. We take a model-based approach to achieve sample efficiency by sequentially updating and improving the dynamics model estimates based on the previously observed transitions. At the beginning of each episode $n$, the representative agent updates its model based on $\cup_{i=1}^{n-1}\mathcal{D}_{i}$ where $\mathcal{D}_{i}=\{(z_{i,t},s_{i,t+1})\}_{t=0}^{T-1}$ and $z_{i,t}=(s_{i,t},\mu_{i,t},a_{i,t})$ is the set of observations in episode $i$ for $i=1,...,n-1$, i.e., up until the beginning of episode $n$. We estimate the mean vector $\bm{m}_{n-1}:\mathcal{Z}\rightarrow\mathcal{S}$ and covariance matrix $\bm{\Sigma}_{n-1}:\mathcal{Z}\rightarrow\mathbb{R}^{p\times p}$ from the set of collected trajectories $\cup_{i=1}^{n-1}\mathcal{D}_{i}$, and denote model’s confidence with $\bm{\sigma}_{n-1}(z)=\text{diag}(\bm{\Sigma}_{n-1}(z))$. We assume that the statistical model is calibrated, meaning that at the beginning of every episode, the agent has high probability confidence bounds around unknown dynamics. The following assumptions are consistent with [50, 12, 15, 47, 43] and other literature which aims to exclude extreme functionals from consideration.
Assumption 4 (Calibrated model).
Let $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ be the mean and covariance functions of the statistical model of $f$ conditioned on $n$ observed episodes. For the confidence function $\bm{\sigma}_{n-1}(\cdot)$, there exists a non-decreasing, strictly positive sequence $\{\beta_{n}\}_{n\geq 0}$ such that for $\delta>0$ with probability at least $1-\delta$, we have jointly for all $n\geq 1$ and $z\in\mathcal{Z}$ that $|f(z)-\bm{m}_{n-1}(z)|\leq\beta_{n-1}\bm{\sigma}_{n-1}(z)$ elementwise.
Assumption 5 (Estimated confidence Lipschitz continuity).
The confidence function $\bm{\sigma}_{n}(\cdot)$ is $L_{\sigma}$-Lipschitz-continuous for all $n\geq 0$, i.e., $\|\bm{\sigma}_{n}(z)-\bm{\sigma}_{n}(z^{\prime})\|_{2}\leq L_{\sigma}d(z,z^{\prime})$.
Since the true transition dynamics model is unknown in Equation 4c and Equation 4d, at the beginning of every episode $n$, the representative agent can only construct the confidence set of dynamics functions $\mathcal{F}_{n-1}$ with $\bm{m}_{n-1}(\cdot)$ and $\bm{\sigma}_{n-1}(\cdot)$ estimated based on the observations up until the end of the previous episode $n-1$, i.e.,
$\mathcal{F}_{n-1}=\Big{\{}\tilde{f}:|\tilde{f}(z)-\bm{m}_{n-1}(z)|\leq\beta_{n-1}\bm{\sigma}_{n-1}(z),\text{ elementwise and }\forall z\in\mathcal{Z}\Big{\}}.$ (5)
The crucial challenge in Safe Mean-Field Reinforcement Learning is that the representative agent can only select dynamics $\tilde{f}\in\mathcal{F}_{n-1}$ at the beginning of the episode $n$ and use it instead of true dynamics $f$ when solving Section 3.1 to find an optimal policy profile $\bm{\pi}_{n}^{*}$. The resulting mean-field distributions $\{\tilde{\mu}_{t}\}_{t=1}^{T}$ are then different from $\{\mu_{t}\}_{t=1}^{T}$ (i.e., the ones that correspond to the true dynamics model), and hence the constraint Equation 3 guarantees only the safety under the dynamics $\tilde{f}$, i.e., $h_{C}(\tilde{\mu}_{t})\geq 0$. In contrast, the original environment constraint $h_{C}(\mu_{t})\geq 0$ might be violated, resulting in unsafe mean-field distributions under the true dynamics $f$.
Next, we demonstrate how to modify the constraint Equation 4e for the optimization problem Section 3.1 when a dynamics function $\tilde{f}$ is used from the confidence set $\mathcal{F}_{n-1}$ such that the mean-field distributions induced by the resulting policy profile do not violate the original constraint under the true dynamics $f$. First, we require the following property for any safety function $h(\cdot)$:
Assumption 6 (Safety Lipschitz continuity).
The safety function $h(\cdot)$ is $L_{h}$-Lipschitz-continuous, i.e., $|h(\mu)-h(\mu^{\prime})|\leq L_{h}W_{1}(\mu,\mu^{\prime}).$
The previous assumption implies that the safety constraint $h_{C}(\cdot)$ exhibits analogous property. The following lemma shows that we can ensure safety under the true dynamics $f$ by having a tighter constraint under any dynamics $\tilde{f}$ selected from $\mathcal{F}_{n-1}$.
Lemma 1.
Given a fixed policy profile $\bm{\pi}_{n}$, a safety function $h(\cdot)$ satisfying Assumption 6 and a safety threshold $C\in\mathbb{R}$, we have in episode $n$ for all steps $t$:
$|h_{C}(\tilde{\mu}_{n,t})-h_{C}(\mu_{n,t})|\leq L_{h}C_{n,t},$
where $C_{n,t}$ is an arbitrary constant that satisfies $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$.
Proof.
For arbitrary $\tilde{\mu}_{n,t},\mu_{n,t}\in\mathcal{P}(\mathcal{S})$ we have $|h_{C}(\tilde{\mu}_{n,t})-h_{C}(\mu_{n,t})|=|h(\tilde{\mu}_{n,t})-h(\mu_{n,t})|\leq L_{h}W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})\leq L_{h}C_{n,t}$, where the first equality follows from the definition of $h_{C}(\cdot)$, the first inequality follows from Assumption 6 and the second inequality comes from $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$.∎
Crucially, using Lemma 1 we can formulate a safety constraint for the optimization under dynamics $\tilde{f}$ that ensure that the constraint under the true dynamics $f$ is not violated.
Corollary 1.
For every episode $n$ and every step $t$, $h_{C}(\tilde{\mu}_{n,t})\geq L_{h}C_{n,t}$ implies $h_{C}(\mu_{n,t})\geq 0$ guaranteeing the safety of the original system.
Proof.
The corollary follows directly from Lemma 1 and the triangle inequality:
$L_{h}C_{n,t}\leq h_{C}(\tilde{\mu}_{n,t})\leq|h_{C}(\tilde{\mu}_{n,t})-h_{C}(\mu_{n,t})|+h_{C}(\mu_{n,t})\leq L_{h}C_{n,t}+h_{C}(\mu_{n,t}),$
where the second inequality follows from the triangle inequality and the third from Lemma 1. The corollary then follows by subtracting the positive constant $L_{h}C_{n,t}$ from both sides. ∎
Then, $C_{n,t}$ for $t=1,\dots,T$ become parameters of the optimization problem (as defined in Section 4) that the representative agent faces at the beginning of episode $n$. However, choosing the appropriate values that comply with the condition $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ is not trivial since $\mu_{n,t}$ depends on the unknown true dynamics of the system. Nevertheless, note that computing $C_{n,0}$ at the initial step $t=0$ is not necessary because the inequality is always guaranteed due to the initialization $\tilde{\mu}_{n,0}=\mu_{n,0}$ for every episode $n$. In Appendix A, we demonstrate how to efficiently upper bound $W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ and obtain $C_{n,t}$ using the Lipschitz constants of the system and the statistical model’s epistemic uncertainty. In particular, $C_{n,t}$ approaches zero, and $h_{C}(\tilde{\mu}_{n,t})\geq L_{h}C_{n,t}$ reduces to the constraint Equation 4e as $\bm{\sigma}_{n-1}(\cdot)$ shrinks due to the increasing number of observations available to estimate the true dynamics.
3.3 Examples of Safety Constraints
We can model multiple classes of safety constraints $h_{C}(\cdot)\geq 0$ that naturally appear in real-world applications such as vehicle repositioning, traffic flow, congestion control, and others. We provide concrete examples of safety functions in Appendix B together with proofs that they satisfy Assumption 6.
Entropic safety. Entropic constraints can be used in multi-agent systems to prevent overcrowding by promoting spatial diversity and preventing excessive clustering. Incorporating an entropic term in the decision-making process encourages the central controller to distribute the agents evenly within the state space. This can be particularly useful in applications that include crowd behavior, such as a swarm of drones or a fleet of vehicles. In such scenarios, we define safety by imposing a hard threshold $C>0$ on the entropy of the mean-field distribution $\mu$:
$h_{C}(\mu)=-\int_{\mathcal{S}}\log\mu(s)\mu(ds)-C.$
Distribution similarity. Another way to define safety is by preventing $\mu$ from diverging from a prior distribution $\nu_{0}$. The prior can be based on previous studies, expert opinions or regulatory requirements. We can use a penalty function that quantifies the allowed dissimilarity between the two distributions:
$h_{C}(\mu;\nu_{0})$ $=C-D(\mu,\nu_{0}),$
with $C>0$ and where the distance function between probability measures $D:\mathcal{P}(\mathcal{S})\times\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{+}$ depends on the problem at hand.
4 Safe-$\text{M}^{3}$-UCRL
In this section, we introduce a model-based approach for the Safe Mean-Field Reinforcement Learning problem that combines the safety guarantees in Corollary 1 with upper-bound confidence interval optimization. At the beginning of each episode $n$, the representative agent constructs the confidence set of dynamics functions $\mathcal{F}_{n-1}$ (see Equation 5) given the calibrated model and previously observed data and selects a safe optimistic policy profile $\bm{\pi}_{n}^{*}$ to obtain the highest value function within $\mathcal{F}_{n-1}$ while satisfying the safety constraint derived in Corollary 1. In particular, the optimal policy profile $\bm{\pi}^{*}$ from Section 3.1 is approximated at the episode $n$ by:
$\bm{\pi}_{n}^{*}$ $=\operatorname*{arg\,max}_{\bm{\pi}_{n}\in\Pi}\max_{\tilde{f}_{n-1}\in\mathcal{F}_{n-1}}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})\Big{\rvert}\tilde{\mu}_{n,0}=\mu_{0}\Bigg{]}$ (6a)
s.t. $\quad\tilde{a}_{n,t}=\pi_{n,t}(\tilde{s}_{n,t},\tilde{\mu}_{n,t})$ (6b)
$\tilde{s}_{n,t+1}=\tilde{f}_{n-1}(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})+\varepsilon_{n,t}$ (6c)
$\tilde{\mu}_{n,t+1}=U(\tilde{\mu}_{n,t},\pi_{n,t},\tilde{f}_{n-1})$ (6d)
$h_{C}(\tilde{\mu}_{n,t+1})\geq L_{h}C_{n,t+1}.$ (6e)
Section 4 optimizes over the function space $\mathcal{F}_{n-1}$ which is usually intractable even in bandit settings [16]. Additionally, it must comply with the safety constraint Equation 6e, further complicating the optimization. We utilize the hallucinated control reparametrization and the log-barrier method to alleviate these issues. After the reformulation of the problem, model-free or model-based mean-field optimization algorithms can be applied to find policy profile $\bm{\pi}_{n}^{*}$ at the beginning of episode $n$.
We use the established approach known as Hallucinated Upper Confidence Reinforcement Learning (H-UCRL) [39, 15, 43] and introduce an auxiliary function $\eta:\mathcal{Z}\rightarrow[-1,1]^{p}$, where $p$ is the dimensionality of the state space $\mathcal{S}$, to define hallucinated dynamics model:
$\tilde{f}_{n-1}(z)=\bm{m}_{n-1}(z)+\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z).$ (7)
Notice that $\tilde{f}_{n-1}$ is calibrated for any $\eta(\cdot)$ under Assumption 4, i.e., $\tilde{f}_{n-1}\in\mathcal{F}_{n-1}$. Assumption 4 further guarantees that every function $\tilde{f}_{n-1}$ can be expressed in the auxiliary form Equation 7:
$\forall\tilde{f}_{n-1}\in\mathcal{F}_{n-1}\;\exists\eta:\mathcal{Z}\rightarrow[-1,1]^{p}\;\text{such that}\;\tilde{f}_{n-1}(z)=\bm{m}_{n-1}(z)+\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),\;\forall z\in\mathcal{Z}.$
Thus, the optimization problem in Section 4 can be expressed as an optimization over the set of admissible policies $\Pi$ and auxiliary function $\eta(\cdot)$ (see Section C.2 for further details). Note that $\eta(z)=\eta(s,\mu,\pi(s,\mu))=\eta(s,\mu)$ for a fixed individual policy $\pi$. This turns $\eta(\cdot)$ into a policy that exerts hallucinated control over the epistemic uncertainty of the confidence set of dynamics functions $\mathcal{F}_{n-1}$ [15]. Furthermore, Equation 7 allows us to optimize over parametrizable functions (e.g., neural networks) $\bm{\pi},\eta(\cdot),\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ using gradient ascent.
We also introduce the safety constraint to the objective using the log-barrier method [55]. This method restricts the domain on which the objective function is defined only to values that satisfy the constraint Equation 6e, hence, turning Section 4 to an unconstrained optimization problem. Combining these two methods yields the following optimization problem:
$\bm{\pi}_{n}^{*}=\operatorname*{arg\,max}_{\bm{\pi}_{n}\in\Pi}$ $\max_{\eta(\cdot)\in[-1,1]^{p}}\mathbb{E}\Bigg{[}\sum_{t=0}^{T-1}r(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})+\lambda\log(h_{C}(\tilde{\mu}_{n,t+1})-L_{h}C_{n,t+1})\Big{\rvert}\tilde{\mu}_{n,0}=\mu_{0}\Bigg{]}$ (8a)
[…]
$\tilde{f}_{n-1}(\tilde{z})$ $=\bm{m}_{n-1}(\tilde{z})+\beta_{n-1}\bm{\Sigma}_{n-1}(\tilde{z})\eta(\tilde{z})$ (8c)
$\tilde{s}_{n,t+1}$ $=\tilde{f}_{n-1}(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})+\varepsilon_{n,t}$ (8d)
$\tilde{\mu}_{n,t+1}$ $=U(\tilde{\mu}_{n,t},\pi_{n,t},\tilde{f}_{n-1}),$ (8e)
with $\lambda>0$ being a tuneable hyperparameter used to balance between the reward and the safety constraint. Provided that the set of safe mean-field distributions (including the initial distribution $\mu_{0}$ that we initialize ourselves) is not empty, $\bm{\pi}_{n}^{*}$ is guaranteed not to violate the safety constraint during the policy rollout in episode $n$.
Remark 1.
Note that Section 4 can also be used under the assumption of known true dynamics by setting $\bm{m}_{n-1}(\cdot)=f(\cdot)$, $\bm{\Sigma}_{n-1}(\cdot)=0$ and $L_{h}C_{n,t}=0$, hence, recovering the original constraint $h_{C}(\tilde{\mu}_{n,t})\geq 0$ from Equation 3. In Section 5, we utilize this useful algorithm property to construct the upper bound for the reward obtained with unknown dynamics.
Algorithm 1 Model-based Learning Protocol in Safe-$\text{M}^{3}$-UCRL
1:Set of admissible policy profiles $\Pi$, safety constraint $h_{C}(\cdot)$, calibrated statistical model represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known reward $r(\cdot)$, safety Lipschitz constant $L_{h}$, hyperparameter $\lambda$, number of episodes $N$, number of steps $T$
2:for $n=1,\ldots N$ do
3:     Compute $C_{n,t}$ for $t=1,\ldots,T$
4:     Optimize the objective in Section 4 with respect to the policy profile and hallucinated dynamics
5:     Execute the obtained policy profile and collect the trajectories from the representative agent
6:     Update the confidence set of dynamics $\mathcal{F}_{n-1}$ with the collected data to obtain $\mathcal{F}_{n}$ for the next episode
7:end for
8:$\bm{\pi}_{N}^{*}=(\pi_{N,0}^{*},\ldots,\pi_{N,T-1}^{*})$
We summarize the model-based learning protocol used by Safe-$\text{M}^{3}$-UCRL in Algorithm 1. The first step consists of computing constants $C_{n,t}$ (see Appendix A) introduced in Lemma 1. The second step is optimizing the objective in Section 4. The third and fourth steps collect trajectories from the representative agent and update the calibrated model. While the learning protocol is model-based, the subroutine in Line 3 can be implemented with either model-based or model-free algorithms proposed for the Mean-Field Control due to our reformulation in Section 4. In Section C.3, we introduce modifications of well-known algorithms for optimizing the mean-field setting.
5 Experiments
In this section, we demonstrate the performance of Safe-$\text{M}^{3}$-UCRL on the vehicle repositioning problem of a ride-hailing platform. Since ride-hailing services, such as Uber and Lyft, gained popularity and market share, vehicle repositioning has been a long-standing challenge for these platforms, i.e., moving idle vehicles to areas with high-demand potential. In the competitive environment, the operator significantly increases profit by successfully repositioning idle vehicles to high-demand areas. Nevertheless, there might exist regulations imposed by the country or city that enforce service providers to either guarantee fair service accessibility or restrict the number of vehicles in districts with high traffic density. Such restrictions prevent operators from greedily maximizing the profit and can be encapsulated by some dispersion metric such as entropy. Solving this problem helps prevent prolonged vehicle cruising and extensive passenger waiting times in the demand hotspots, increasing the service provider’s efficiency and reducing its carbon footprint. Existing approaches to vehicle repositioning range from static optimization over a queuing network [56, 7], model predictive control [27], to reinforcement learning [54, 35, 37]. The main advantage of Safe-$\text{M}^{3}$-UCRL is the capability of controlling a large fleet of homogeneous vehicles and enforcing efficient coordination to match the spatiotemporal demand distribution. Additionally, the safety constraint introduced into the model guarantees service accessibility by ensuring idle vehicles are spreading over the study region. Although accessibility has not been widely discussed in the literature on vehicle repositioning, it is expected to be an important fairness constraint when shared mobility services become a prevailing travel mode [48].
Observed demand used as a target distribution $\rho_{0}$
$\text{M}^{3}$-UCRL with known dynamics
Safe-$\text{M}^{3}$-UCRL with known dynamics
Safe-$\text{M}^{3}$-UCRL with unknown dynamics
Figure 2: Safe-$\text{M}^{3}$-UCRL performance in Shenzhen. We show the final mean-field distributions $\mu_{T}$ for $T=12$ and the entropic safety constraint $h_{C}(\cdot)$ (Section 3.3). We set the safety threshold as $C=0.85\log(k^{2})$ with $k=25$ which represents 85% of the maximum Shannon entropy. The observed demand distribution $\rho_{0}$ in (a) horizontally spreads across the city center, which amounts to 67% of the maximum entropy, while residential areas in the northwest and especially northeast show only glimpses of demand. Still, the service providers might be obligated to guarantee coverage in these areas. (b) shows that $\text{M}^{3}$-UCRL , the benchmark model without safety constraint, closely matches the target distribution, resulting in safety constraint violation as shown in Figure 2(b). (c) shows that our model with known dynamics achieves coverage in all residential areas while keeping the majority of taxis in demand hotspots. (d) shows the competitive performance of our model even when faced with unknown dynamics.
Modeling. Ride-hailing operations can be modeled as sequential decision-making, which consists of passenger trips followed by repositioning trips operated by a central controller as illustrated in Figure 1. We assume that the controller has access to the locations of taxis in its fleet and that real-time communication of repositioning actions to the taxi drivers is possible via electronic devices. Nevertheless, since the fleet is operating in a noisy traffic environment, repositioning usually cannot be executed perfectly. We assume that taxis can move freely within the area of our interest, which is represented by a two-dimensional unit square, i.e., the state-space $\mathcal{S}=[0,1]^{2}$, and repositioning actions are taken from $\mathcal{A}=[-1,1]^{2}$. The objective of our model is to satisfy the demand in the central district of Shenzhen while providing service accessibility in the wider city center. We restrict our modeling horizon to three evening peak hours which are discretized in fifteen-minute operational intervals, i.e., $T=12$, and each episode $n$ represents one day (for more details, see Appendix D). We implicitly model the profit by the negative of the Kullback-Leibler divergence between the taxis’ distribution $\mu_{n,t}$ and demand for taxi service denoted as $\rho_{0}$, i.e., $r(s_{n,t},\mu_{n,t},a_{n,t})=-D_{KL}(\rho_{0}||\mu_{n,t})$. In particular, the demand distribution $\rho_{0}\in\mathcal{P}(\mathcal{S})$ represents a probability of a trip originating in the infinitesimal neighborhood of state $s\in\mathcal{S}$ during peak hours (see Figure 1(a)). We estimate the demand distribution $\rho_{0}$ from the taxi trajectories collected in Shenzhen, China, in 2016 and keep it fixed for every step $t$. If the passenger’s trip originates at $s\in\mathcal{S}$ the likelihood of its destinations is defined by the mapping $\Phi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ which we fit from the trip trajectories (see Appendix D). We use $\Phi(\cdot)$ to define sequential transitions by first executing passenger trips followed by vehicle repositioning. Formally, the unknown system dynamics are given by $f(s_{n,t},\mu_{n,t},a_{n,t})=\text{clip}(s^{\Phi}_{n,t}+a_{n,t},0,1)+\varepsilon_{n,t}$, where $s^{\Phi}_{n,t}=\text{sample}(\Phi(s_{n,t}))$ and $\varepsilon_{n,t}\sim\text{TN}(0,\sigma^{2}I_{2})$ is a Gaussian with a known variance $\sigma^{2}$ truncated at the borders of $\mathcal{S}$ and $I_{2}$ is the $2\times 2$ unit matrix. We use entropic safety constraint $h_{C}(\mu_{n,t})=-\int_{\mathcal{S}}\log(\mu_{n,t}(s))\mu_{n,t}(ds)-C$ to enforce the service accessibility across all residential areas (see Figures 1(c) to 1(d)). Therefore, the optimization objective in Section 4 trades off between satisfying the demand $\rho_{0}$ and adhering to accessibility constraint imposed by $h_{C}(\cdot)$.
We use neural network to parametrize the policy profile $\bm{\pi}_{n}=(\pi_{n,0},\ldots,\pi_{n,T-1})$, for every episode $n$, during the optimization of Section 4. The optimization is done by a mean-field back-propagation through time (MF-BPTT) (see Section C.3.1). In our experiments, a single neural network shows enough predictive power to represent the whole policy profile, but using $T$ networks, one for each individual policy $\pi_{n,t}$ is a natural extension. We use a Probabilistic Ensemble to represent a statistical model of system dynamics, which is elaborated in Section C.1. We represent the mean-field distribution by discretizing the state space into $k\times k$ grid with $k=25$ and assigning the probability of the representative agent residing within each cell. The initial mean-field distributions $\mu_{n,0}$ follow the uniform distribution to maximize the Shannon entropy, ensuring the safety at $t=0$, i.e., $h_{C}(\mu_{n,0})\geq 0$ for every episode $n$. We set the safety threshold $C$ as a percentage $p$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$.
(a) Objective from Equation 8a achieved
over 200 training episodes
(b) Constraint violations comparison
against the benchmark
Figure 3: Safe-$\text{M}^{3}$-UCRL training and evaluation. We train 10 randomly initialized policy profiles and show that: In (a) the performance of policy profiles during the exploration is quite volatile, but around episode 80, policy profiles become confident about the underlying dynamics and quickly converge to a solution as if the dynamics were known; In (b) our policy profiles deployed in the true environment never violate the entropic safety constraint $h_{C}(\cdot)\geq 0$ for $C=0.85\log(k^{2})$ with $k=25$, while the benchmark model $\text{M}^{3}$-UCRL trained without the safety constraint starts displaying unsafe behavior after only two steps.
Results. The target distribution $\rho_{0}$ in Figure 1(a) has a relatively high entropy due to a wide horizontal spread of $67\%$ of the maximum entropy. To achieve vertical spread, we require additional 18 percentage points of entropy as a safety constraint. Concretely, we use $p=85\%$ to set the hard constraint $C=p\log(k^{2})$ and proceed by optimizing the policy profile in Section 4. Due to the lack of an analytical solution for Section 4, we use a policy profile trained with known dynamics as a benchmark. We show that $\bm{\pi}_{n}^{*}$ converges to this policy profile in $n=80$ episodes. Figure 2(a) shows two phases of the learning process. During the first 60 episodes, the performance is volatile, and then the confidence bound shrinks, i.e., the random exploration reduces. Once the epistemic uncertainty around the true dynamics is tight, the model exploits it rapidly by episode 80.
By comparing Safe-$\text{M}^{3}$-UCRL to the benchmark $\text{M}^{3}$-UCRL introduced in [43], which does not impose any safety constraint, we empirically show that Safe-$\text{M}^{3}$-UCRL satisfies the safety constraint during the entire execution. In contrast, $\text{M}^{3}$-UCRL quickly violates it (see Figure 2(b)). In Figure 2, we use a city map of Shenzhen to show that Safe-$\text{M}^{3}$-UCRL improves service accessibility in low-demand areas. Figure 1(b) shows that $\text{M}^{3}$-UCRL with known dynamics learns how to satisfy the demand $\rho_{0}$ effectively, but with the discussed drawback of violating safety constraint. Figure 1(c) shows that Safe-$\text{M}^{3}$-UCRL with known dynamics improves on that by satisfying safety constraint by distributing taxis to residential areas in the northwest and northeast. Finally, Figure 1(d) emphasizes the capability of Safe-$\text{M}^{3}$-UCRL to learn complex system dynamics while the policy profile $\bm{\pi}_{n}^{*}$ simultaneously converges towards the results achieved with known dynamics with the number of episodes $n$ passed.
In Appendix D, we provide the details on the parameters used during the training and exhaustive performance analysis. The results in this section are generated assuming the infinite regime. At the same time, in Appendix D, we also discuss the performance when the policy profile $\bm{\pi}_{n}^{*}$ is applied to the finite number of individual agents. The code we used to train and evaluate Safe-$\text{M}^{3}$-UCRL is available at https://github.com/mjusup1501/safe-m3-ucrl.
6 Conclusion
We present a novel formulation of a mean-field model-based reinforcement learning problem incorporating safety constraints. Safe-$\text{M}^{3}$-UCRL addresses this problem by leveraging epistemic uncertainty in the unknown transition dynamics model and employing a log-barrier approach to ensure conservative satisfaction of the constraints. Through experimental evaluations, we showcase the effectiveness of our algorithm in matching the demand distribution in a shared mobility service while consistently upholding service accessibility. In the future, we believe that integrating safety considerations in intelligent multi-agent systems will have a crucial impact on various applications, such as firefighting robots and drone/robot search-and-rescue operations in complex and confined spaces (e.g., mining accidents or tunnel collapses).
Acknowledgments and Disclosure of Funding
The authors would like to thank Mojmír Mutný for the fruitful discussions during the course of this work. Matej Jusup acknowledges support from the Swiss National Science Foundation under the research project DADA/181210. This publication was made possible by an ETH AI Center doctoral fellowship to Barna Pasztor. Francesco Corman acknowledges Grant 2023-FS-331 for Research in the area of Public Transport. The authors would also like to thank the Shenzhen Urban Transport Planning Center for collecting and sharing the data used in this work.
Appendix
Safe Model-Based Multi-Agent
Mean-Field Reinforcement Learning
Appendix A Relationship Between True and Statistical Mean-Field Distributions
In this section, we describe the procedure for computing constants $C_{n,t}$ (defined in Section 4) at every episode $n=1,\ldots,N$ and step $t=1,\ldots,T$. These are necessary for establishing the connection between the mean-field distributions $\mu_{n,t}$ in the original system with known transition dynamics $f$ (Section 3.1) and the mean-field distributions $\tilde{\mu}_{n,t}$ in the system induced by a calibrated statistical model $\tilde{f}_{n-1}$ (Section 4, Section 3.2 and Assumption 4). In particular, Corollary 1 shows that the safety of the original system $h_{C}(\mu_{n,t})$ is guaranteed if $h_{C}(\tilde{\mu}_{n,t})\geq L_{h}C_{n,t}$ for a safety constant $C_{n,t}\geq W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ and Lipschitz constant $L_{h}$ introduced in Assumption 6. We use the following result from [43, Lemma 5] to make a connection between the mean-field distributions in the true dynamics and statistical model dynamics:
Lemma 2.
Under Assumptions 1 to 5 and assuming that the event in Assumption 4 holds true, for episodes $n=1,\ldots,N$, steps $t=1,\ldots,T$ and fixed policy profile $\bm{\pi}_{n}=(\pi_{n,0},\ldots,\pi_{n,T-1})$, we have:
$W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})\leq 2\beta_{n-1}\overline{L}_{n-1}^{t-1}\sum_{i=0}^{t-1}\int_{\mathcal{S}}\|\bm{\sigma}_{n-1}(s,\mu_{n,i},\pi_{n,t}(s,\mu_{n,i}))\|_{2}\mu_{n,i}(ds),$
where $\overline{L}_{n-1}=1+2(1+L_{\pi})(L_{f}+2\beta_{n-1}L_{\sigma})$.
Let $K_{n,t}:=2\beta_{n-1}\overline{L}_{n-1}^{t-1}$ and $z=(s,\mu,\pi_{n,t}(s,\mu))$. Then we have:
$W_{1}(\tilde{\mu}_{n,t},\mu_{n,t})$ $\leq K_{n,t}\sum_{i=0}^{t-1}\int_{\mathcal{S}}\|\bm{\sigma}_{n-1}(s,\mu_{n,i},\pi_{n,t}(s,\mu_{n,i}))\|_{2}\mu_{n,i}(ds)$
$\leq tK_{n,t}\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2},$
for $t=1,\ldots,T$, where the first inequality is due to Lemma 2 while the second one follows since maximum upper bounds expectation. Hence, we can set $C_{n,t}$ as $tK_{n,t}\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2}$, and calculate $\max_{z\in\mathcal{Z}}\|\bm{\sigma}_{n-1}(z)\|_{2}$ once at the beginning of each episode $n$. Notice that our learning protocol (Algorithm 1) does not require computing $C_{n,0}$ at the initial step $t=0$ since the mean-field distributions share the initial state, i.e., $\tilde{\mu}_{n,0}=\mu_{n,0}$. It is worth noting that in certain models, such as Gaussian Processes, this upper bound provides a meaningful interpretation. Specifically, $\bm{\sigma}_{n-1}(z)$ represents the epistemic uncertainty of the model, which tends to decrease monotonically as more data is observed.
Appendix B Examples of Safety Constraints
In this section, we show some important classes of safety constraints $h_{C}(\cdot)$ satisfying Assumption 6.
B.1 Entropic Safety
Entropic safety serves the purpose of controlling the dispersion of the mean-field distribution, which is useful in many applications, such as vehicle repositioning (see Section 5). A natural way of defining entropic safety is via differential entropy, but in general, the differential entropy is not Lipschitz continuous due to the unboundedness of the natural logarithm. Nevertheless, the issue can be easily circumvented by considering $\varepsilon$-smoothed differential entropy $H^{\varepsilon}:\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{+}$. For $\varepsilon>0$ and $C>0$ we define $\varepsilon$-smoothed differential entropy and associated entropic safety constraint $H_{C}^{\varepsilon}(\cdot)$ as:
$H^{\varepsilon}(\mu)$ $=-\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)$
$H^{\varepsilon}_{C}(\mu)$ $=H^{\varepsilon}(\mu)-C$
To show that $H^{\varepsilon}_{C}(\cdot)$ satisfies Assumption 6 let $h_{C}(\cdot):=H^{\varepsilon}_{C}(\cdot)$ and assume that $\mathcal{S}\subset\mathbb{R}^{p}$ is a compact set. First, note that $f(x)=\log(x+\varepsilon)$ is $\frac{1}{\varepsilon}$-Lipschitz continuous for $\varepsilon>0$, i.e., $\varepsilon f(x)$ is 1-Lipschitz. Second, for every $S\subseteq\mathcal{S}$ and $L$-Lipschitz function $f:\mathcal{S}\rightarrow\mathbb{R}$, a function $g(S)=\int_{\mathcal{S}}f(s)\mu(ds)$ is $L$-Lipschitz due to the boundedness of $f$. The following derivation shows that $H_{C}^{\varepsilon}(\cdot)$ is $\frac{1}{\varepsilon}$-Lipschitz continuous:
$|h_{C}(\mu)-h_{C}(\mu^{\prime})|$ $=\left|H^{\varepsilon}_{C}(\mu)-H^{\varepsilon}_{C}(\mu^{\prime})\right|$
$=\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)\mu^{\prime}(ds)-\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)\right|$
$=\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)(\mu^{\prime}-\mu)(ds)-\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)+\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)\mu(ds)\right|$
$\leq\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)(\mu^{\prime}-\mu)(ds)\right|+\left|\int_{\mathcal{S}}\log(\mu(s)+\varepsilon)\mu(ds)-\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)\mu(ds)\right|$
$\leq\left|\int_{\mathcal{S}}\log(\mu^{\prime}(s)+\varepsilon)(\mu^{\prime}-\mu)(ds)\right|+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime})$
$\leq M\left|\int_{\mathcal{S}}(\mu^{\prime}-\mu)(ds)\right|+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime})$
$=\underbrace{M(\mu^{\prime}-\mu)(\mathcal{S})}_{\overline{M}}+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime})$
$=\overline{M}+\frac{1}{\varepsilon}W_{1}(\mu,\mu^{\prime}),$
where the first inequality comes from the triangle inequality, the second inequality comes from the Lipschitz continuity of the $\varepsilon$-smoothed logarithm and integral, the third comes from the upper-boundedness of logarithm on a compact set, and the last equality comes from the fact that the measure of a compact set is finite.
Remark 2.
In Section 5 we use the discrete equivalent of entropic safety, i.e., Shannon entropy, for our experiments. [45, Proposition 8] shows that Shannon entropy is Lipschitz continuous with respect to the scaled Wasserstein 1-distance, i.e., with respect to $\frac{1}{n}W_{1}(\cdot)$ known as Ornstein’s distance.
B.2 Distribution Similarity
We can define safety by preventing the mean-field distribution $\mu$ from diverging from a prior distribution $\nu_{0}$ by quantifying the allowed dissimilarity between the two distributions:
$h_{C}(\mu;\nu_{0})$ $=C-D(\mu,\nu_{0}),$
where $C>0$ and $D:\mathcal{P}(\mathcal{S})\times\mathcal{P}(\mathcal{S})\rightarrow\mathbb{R}_{+}$ is the distance function between probability measures. Commonly used distances are Wasserstein p-distance for $p\geq 1$ and $f$-divergences such as KL-divergence, Hellinger distance, and total variation distance.
A concrete example of a distance function $D(\cdot)$ that satisfies Assumption 6 is Wasserstein 1-distance:
$W_{1}^{C}(\mu,\nu_{0})$ $=C-W_{1}(\mu,\nu_{0})$
$h_{C}(\mu)$ $:=W_{1}^{C}(\mu,\nu_{0}).$
[14] shows the triangle inequality of the Wasserstein p-distance for probability measures on separable metric spaces. [17] shows that Wasserstein 1-distance induces a metric space $(\mathcal{P}(\mathcal{S}),W_{1})$ over probability measures. The result now trivially follows from the reverse triangle inequality:
$|h_{C}(\mu)-h_{C}(\mu^{\prime})|$ $=|W_{1}^{C}(\mu,\nu_{0})-W_{1}^{C}(\mu^{\prime},\nu_{0})|$
$=|W_{1}(\mu,\nu_{0})-W_{1}(\mu^{\prime},\nu_{0})|$
$\leq W_{1}(\mu,\mu^{\prime}).$
Remark 3.
Avoiding risky distributions can be modelled by setting $W_{1}^{C}(\mu,\nu_{0})=W_{1}(\mu,\nu_{0})-C$ with the proof of Assumption 6 being equivalent to the above.
Weighted safety constraints. In applications that require emphasis on certain regions of the state space, we can generalize the above safety constraints by introducing the weight function $w:\mathcal{S}\rightarrow\mathbb{R}_{+}$. We extend Section B.1 to weighted-differential entropy by defining $H^{w,\varepsilon}(\mu):=-\int_{\mathcal{S}}w(s)\log(\mu(s)+\varepsilon)\mu(ds)$ and Section B.2 by considering weighted Wasserstein 1-distance $W_{1}^{w}(\mu,\nu_{0})=\sup_{f:\text{Lip}(f)\leq 1}\int_{\mathcal{S}}w(s)f(s)(\mu-\nu_{0})(ds)$. Here, we use Kantorovic-Rubinstein dual definition of Wasserstein 1-distance $W_{1}(\mu,\nu_{0})=\sup_{f:\text{Lip}(f)\leq 1}\int_{\mathcal{S}}f(s)(\mu-\nu_{0})(ds)$, where $f:\mathcal{S}\rightarrow\mathbb{R}$ is a continuous function and $\text{Lip}(f)$ denotes the minimal Lipschitz constant for $f$.
Appendix C Implementation Details
In this section, we provide additional details on the practical implementation of Safe-$\text{M}^{3}$-UCRL. In particular, we discuss Probabilistic Ensemble model [13, 30] to implement the statistical model from Section 3.2 in Section C.1, the hallucinated control reparametrization from Section 4 in more detail in Section C.2, and optimization methods to solve Section 4 in Section C.3
C.1 Probabilistic Ensemble Model of Transition Dynamics
As discussed in Section 3.2, we take a model-based approach to handling the unknown dynamics $f$. The representative agent learns the Statistical Model (see Section 3.2) of the dynamics from the observed transitions $\cup_{i=1}^{n-1}\mathcal{D}_{i}$ at the beginning of each episode $n$, where $\mathcal{D}_{i}=\{(z_{i,t},s_{i,t+1})\}_{t=0}^{T-1}$ with $z_{i,t}=(s_{i,t},\mu_{i,t},a_{i,t})$. We use Probabilistic Ensemble [13, 30] that consist of $K$ neural networks parametrized by $\theta_{k}$ for $k\in\{1,\ldots,K\}$ (the episode index should be clear from the context so we omit it for the notation simplicity). Each neural network $f_{\theta_{k}}$ returns a mean vector, $\bm{m}_{\theta_{k}}(z)\in\mathcal{S}\subseteq\mathbb{R}^{p}$, and a covariance matrix $\bm{\Sigma}_{\theta_{k}}(z)\in\mathbb{R}^{p\times p}$, that represents the aleatoric uncertainty. We further assume diagonal covariance matrices. These outputs then form Gaussian distributions from which new states are sampled, i.e., $s_{t+1}\sim\mathcal{N}(\bm{m}_{\theta_{k}}(z_{t}),\bm{\Sigma}_{\theta_{k}}(z_{t}))$ for $t=0,\ldots,T-1$. The models are trained with the negative log-likelihood loss function (NLL), $L(\theta)=-\sum_{\mathcal{D}_{1:n-1}}\log\mathbb{P}(s_{t+1}|z_{t})$ as described in [30]. The ensemble means and the aleatoric and epistemic uncertainties are then estimated as follows:
$\bm{m}_{n-1}(\cdot)$ $=\frac{1}{K}\sum_{k=1}^{K}\bm{m}_{\theta_{k}}(\cdot),$
$\bm{\Sigma}_{n-1}^{e}(\cdot)$ $=\frac{1}{K-1}\sum_{k=1}^{K}(\bm{m}_{\theta_{k}}(\cdot)-\bm{m}_{n-1}(\cdot))(\bm{m}_{\theta_{k}}(\cdot)-\bm{m}_{n-1}(\cdot))^{\top},$
$\bm{\Sigma}_{n-1}^{a}(\cdot)$ $=\frac{1}{K}\sum_{k=1}^{K}\bm{\Sigma}_{\theta_{k}}(\cdot),$
where $\bm{m}_{n-1}(\cdot)$ is the ensemble prediction for the mean, while $\bm{\Sigma}_{n-1}^{a}(\cdot)$ and $\bm{\Sigma}_{n-1}^{e}(\cdot)$ denote the aleatoric and epistemic uncertainty estimates, respectively. Note that the epistemic estimate $\bm{\Sigma}_{n-1}^{e}(\cdot)$ is used to construct the calibrated model (see Section 3.2 and Assumption 4).
Even though Gaussian Processes (GPs) are proven to be calibrated under certain regularity assumptions [1, 50], we chose probabilistic ensemble model due to their better practical performance, i.e., scalability to higher dimensions and larger datasets. The disadvantage of the probabilistic ensemble is that, unlike GPs, it does not guarantee the calibrated model (see Assumption 4). Nevertheless, it can be recalibrated using one-step ahead predictions and temperature scaling as shown in [36]. We note that in our experiment (see Section 5), such recalibration was not needed, and the above-defined mean and epistemic uncertainty outputs were sufficiently accurate for training Safe-$\text{M}^{3}$-UCRL.
C.2 Hallucinated Control Implementation Trick
In Section 4, we introduced Safe-$\text{M}^{3}$-UCRL, a model that optimizes over the confidence set of dynamics functions $\mathcal{F}_{n-1}$ and admissible policy profiles $\Pi$ (see Section 4). Unfortunately, optimizing directly over the function space is usually intractable since $\mathcal{F}_{n-1}$ is not convex, in general, [16]. Thus, to make the optimization tractable, we describe a hallucinated control trick, which leads to a practical reformulation (see Section 4). The structure in $\mathcal{F}_{n-1}$ allows us to parametrize the problem and use gradient-based optimization to find a policy profile $\bm{\pi}_{n}^{*}$ at every episode $n$. Namely, we use a mean-field variant of an established approach known as Hallucinated Upper Confidence Reinforcement Learning (H-UCRL) [15, 39, 43]. We introduce an auxiliary function $\eta:\mathcal{Z}\rightarrow[-1,1]^{p}$, where $p$ is the dimensionality of the state space $\mathcal{S}$, to define hallucinated dynamics model:
$\tilde{f}_{n-1}(z)=\bm{m}_{n-1}(z)+\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),$ (9)
where $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ are estimated from the past observations collected until the end of the previous episode $n-1$. Notice that $\tilde{f}_{n-1}$ is calibrated for any $\eta(\cdot)$ under Assumption 4, i.e., $\tilde{f}_{n-1}\in\mathcal{F}_{n-1}$. Assumption 4 further guarantees that every function $\tilde{f}_{n-1}$ can be expressed in the auxiliary form in Equation 9:
$\forall\tilde{f}_{n-1}\in\mathcal{F}_{n-1}\;\exists\eta:\mathcal{Z}\rightarrow[-1,1]^{p}\;\text{such that}\;\tilde{f}_{n-1}(z)=\bm{m}_{n-1}(z)+\beta_{n-1}\bm{\Sigma}_{n-1}(z)\eta(z),\;\forall z\in\mathcal{Z}.$
Furthermore, note that, for a fixed individual policy $\pi$, the auxiliary function $\eta(z)=\eta(s,\mu,\pi(s,\mu))=\eta(s,\mu)$ has the same functional form as the policy $\pi$. This turns $\eta(\cdot)$ into a policy that exerts hallucinated control over the epistemic uncertainty of the confidence set of dynamics functions $\mathcal{F}_{n-1}$ [15]. The reformulation of the optimization problem in Section 4 allows us to optimize over parametrizable functions (e.g., neural networks) $\bm{\pi},\eta(\cdot),\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$ using gradient-based methods on the functions’ parameters. Note that the parametrization of $\eta(\cdot)$ must be sufficiently flexible not to restrict the space of $\tilde{f}_{n-1}$. In Section C.3, we provide several algorithms that can be used to solve this optimization problem.
C.3 Optimization Methods (MF-BPTT & MF-DDPG)
In this section, we describe two algorithms to solve the optimization problem in Section 4. Namely, in Section C.3.1, we outline the key steps to apply the mean-field variant of the Back Propagation Through Time (BPTT) when a differentiable simulator is available and, in Section C.3.2, we describe the mean-field variant of the Deep Deterministic Policy Gradient (DDPG) [34] algorithm.
C.3.1 Mean-Field Back-Propagation-Through-Time (MF-BPTT)
Mean-Field Back-Propagation-Through-Time (MF-BPTT) assumes access to a differentiable simulator that returns a policy rollout given transition and policy functions. In each episode $n$, the representative agent initializes the policy profile $\bm{\pi}_{n}$, the auxiliary function $\eta(\cdot)$, and the estimated dynamics $\tilde{f}_{n-1}$ using the mean vector $\bm{m}_{n-1}(\cdot)$ and covariance matrix $\bm{\Sigma}_{n-1}(\cdot)$ according to Equation 7. Then, the representative agent repeatedly calls the simulator with inputs $\bm{\pi}_{n}$ and $\tilde{f}_{n-1}$ that returns the episode reward defined in Equation 8a as a differentiable object. After each policy rollout, a gradient ascent step is carried out on the parameters of $\bm{\pi}_{n}$ and $\eta(\cdot)$ before calling the simulator again. To simplify the notation, we overload the notation $\bm{\pi}_{n}^{\psi}$ with the combination of the two policy functions $\bm{\pi}_{n}$ and $\eta(\cdot)$, i.e., $\bm{\pi}_{n}^{\psi}=(\bm{\pi}_{n}^{\psi},\eta^{\psi})$ where the superscript $\psi$ represent the parameters of both functions. We outline the described steps in Algorithm 2. During the optimization phase (Line 3 in Algorithm 1), we optimize both functions jointly. However, during the execution (Line 4 in Algorithm 1), we only use the outputs corresponding to the policy profile $\bm{\pi}_{n}$. The main distinction in Algorithm 2 compared to traditional BPTT lies inside the simulator that has to simulate the mean-field distribution $\tilde{\mu}_{n,t}$ for $t=0,\dots,T$ for each parameter update and calculate gradients with respect to this time dependency as well. More details on the implementation used for our experiments reported in Section 5 are provided in Appendix D.
Algorithm 2 Mean-Field Back-Propagation-Through-Time
1:Safety constraint $h_{C}(\cdot)$, calibrated dynamics $\tilde{f}_{n-1}$ represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known reward $r(\cdot)$, constants $C_{n,t}$, safety Lipschitz constant $L_{h}$; hyperparameter $\lambda$, number of epochs $K$, number of rollout steps $T$
2:Initialize $\bm{\pi}^{\psi}=(\pi^{\psi}_{0},\ldots,\pi^{\psi}_{T-1})$
3:for $k=1,\ldots,K$ do
4:     Initialize $\tilde{\mu}_{0}\leftarrow\mu_{0}$, $\tilde{s}_{0}\sim\mu_{0}$
5:     Initialize $r\leftarrow 0$
6:     for $t=0,\ldots,T-1$ do
7:         $\tilde{a}_{t}\leftarrow\pi^{\psi}_{t}(\tilde{s}_{t},\tilde{\mu}_{t})$
8:         $\tilde{s}_{t+1}\leftarrow\tilde{f}_{n-1}(\tilde{s}_{t},\tilde{\mu}_{t},\tilde{a}_{t})+\varepsilon_{t}$
9:         $\tilde{\mu}_{t+1}\leftarrow U(\tilde{\mu}_{t},\pi^{\psi}_{t},\tilde{f}_{n-1})$
10:     end for
11:     Update $\psi$ with a gradient ascent:
12:     $\quad\;\,\nabla_{\psi}\sum_{t=0}^{T-1}r(\tilde{s}_{t},\tilde{\mu}_{t},\tilde{a}_{t})+\lambda\log(h_{C}(\tilde{\mu}_{t+1})-L_{h}C_{n,t+1})$
13:end for
14:$\bm{\pi}_{n}^{\psi}\leftarrow\bm{\pi}^{\psi}$
C.3.2 Mean-Field Deep Deterministic Policy Gradient (MF-DDPG)
In this section, we adopt the Deep Deterministic Policy Gradient (DDPG) algorithm [34] to the mean-field control problem. DDPG is a model-free actor-critic algorithm, hence, it can optimize Section 4 without the assumption of a differentiable simulator. However, it can not be applied directly to the problem because the Q-value for $\tilde{z}_{n,t}=(\tilde{s}_{n,t},\tilde{\mu}_{n,t},\tilde{a}_{n,t})$ is ambiguous. The important insight here is that the value of a certain state $\tilde{s}_{n,t}$ of the environment reflects the whole population, i.e., the expected reward over the remainder of an episode for a given $\tilde{\mu}_{n,t}$ if every agent in every state $\tilde{s}_{n,t}$ chooses actions following $\pi_{n,t}$. In essence, the Q-value is a function of $\tilde{\mu}_{n,t}$ and $\pi_{n,t}$ and not $\tilde{z}_{n,t}$.
To overcome this issue, we introduce the lifted mean-field Markov decision process (MF-MDP) similarly to [9, 22, 23, 41, 43]. First, we rewrite the reward function as a function of the mean-field distribution and the policy, i.e.,
$\tilde{r}(\tilde{\mu}_{n,t},\pi_{n,t})=\int_{\mathcal{S}}r(s,\tilde{\mu}_{n,t},\pi_{n,t}(s,\tilde{\mu}_{n,t}))\tilde{\mu}(ds).$
To simplify the notation, we overload the notation $\bm{\pi}_{n}^{\psi}=(\bm{\pi}_{n}^{\psi},\eta^{\psi})$ as described in Section C.3.1. Then, we restate Section 4 as:
$\psi^{*}$ $=\operatorname*{arg\,max}_{\psi}\sum_{t=0}^{T-1}\tilde{r}(\tilde{\mu}_{n,t},\pi_{n,t}^{\psi})+\lambda\log(h_{C}(\tilde{\mu}_{n,t+1})-L_{h}C_{n,t+1})$ (10a)
[…]
$\tilde{\mu}_{n,t+1}$ $=U(\tilde{\mu}_{n,t},\pi_{n,t}^{\psi},\tilde{f}_{n-1}),$ (10c)
where $\tilde{\mu}_{n,0}=\mu_{0}$. The MF-MDP formulation in Section C.3.2 turns the Mean-Field Control problem in Section 4 into a Markov Decision Process on the state space of $\mathcal{P}(\mathcal{S})$ and action space $\{\pi:\mathcal{S}\times\mathcal{P}(\mathcal{S})\to\mathcal{A}\}$ with deterministic transition function $U(\cdot)$. We define the Q-value as follows:
$Q_{n,t}(\tilde{\mu}_{n,t},\bm{\pi}_{n}^{\psi})=\sum_{j=t}^{T-1}\tilde{r}(\tilde{\mu}_{n,t},\pi_{n,t}^{\psi}).$
The Q-function above is parameterized by $\theta$ and denoted as $Q^{\theta}_{n,t}$ for episode $n$. The DDPG algorithm can now be stated for the lifted MF-MDP problem as outlined in Algorithm 3. The main learning loop consists of alternating updates to the policy $\bm{\pi}^{\psi}_{n}$ and the critic $Q^{\theta}_{n}$. The most recent version of the policy is executed in the environment to collect more transitions into the replay buffer.
Notice that Algorithm 3 uses a model-free approach to optimize the objective, which makes the exploration of particular importance. In comparison to traditional MDPs, MF-MDPs usually have a very constrained set of highly rewarding distributions, and most distributions offer poor rewards, which makes the exploration even more important. This is further complicated by what randomized actions imply in this scenario. In a traditional MDP, we can often assume, for instance, that executing random actions for a fixed number of initial steps would help in finding diversity in the reward space. This is not the case in MF-MDPs – depending on the granularity of the discretization that we use to represent probability distributions, we can expect the mean-field distribution to stay fairly stable. Similarly to [9], we might alleviate this issue by Gaussian mean-field initialization in each episode. Note, however, that this may not necessarily be appropriate in safety-constrained settings, as the initial mean-field distribution is expected to be safe. On top of that, we can add exploration noise to the actions obtained via the policy. Alternatively, we could add noise to the parameters of the policy network for a more consistent approach as in [44].
Algorithm 3 Mean-Field Deep Deterministic Policy Gradient
1:Safety constraint $h_{C}(\cdot)$, calibrated dynamics $\tilde{f}_{n-1}$ represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known expected reward $\hat{r}(\cdot)$, constants $C_{n,t}$; safety Lipschitz constant $L_{h}$, hyperparameter $\lambda$, number of epochs $K$, number of rollout steps $T$, mini-batch size $B$, learning rate $\alpha$
2:Initialize $\bm{\pi}^{\psi}=(\pi^{\psi}_{0},\ldots,\pi^{\psi}_{T-1})$
3:Initialize $Q^{\theta}=(Q^{\theta}_{0},\ldots,Q^{\theta}_{T-1})$
4:Initialize $\theta^{\prime}\leftarrow\theta$, $\psi^{\prime}\leftarrow\psi$
5:Initialize replay buffer $R\leftarrow\emptyset$
6:for $k=1,\ldots,K$ do
7:     Initialize $\tilde{\mu}_{0}\leftarrow\mu_{0}$
8:     for $t=0,\ldots,T-1$ do
9:         $\tilde{\mu}_{t+1}\leftarrow U(\tilde{\mu}_{t},\pi^{\psi}_{t},\tilde{f}_{n-1})$
10:         if $t=0$ then
11:              $c_{t}\leftarrow\hat{r}(\tilde{\mu}_{t},\pi^{\psi}_{t})$
12:         end if
13:         if $11\leq t\leq T-$ 1 then
14:              $c_{t}\leftarrow\hat{r}(\tilde{\mu}_{t},\pi^{\psi}_{t})+\lambda\log(h_{C}(\tilde{\mu}_{t})-L_{h}C_{n,t})$
15:         end if
16:         if $t=T$ then
17:              $c_{t}\leftarrow\lambda\log(h_{C}(\tilde{\mu}_{t})-L_{h}C_{n,t})$
18:         end if
19:         $R\leftarrow R\cup\{(\tilde{\mu}_{t},c_{t},\tilde{\mu}_{t+1})\}$
20:         Sample a mini-batch of $B$ random transitions $\{(\tilde{\mu}_{i},c_{i},\tilde{\mu}_{i+1})\}_{i=1}^{B}$ from $R$
21:         for $i=1,\ldots,B$ do
22:              $q_{i}\leftarrow c_{i}+\gamma Q^{\theta^{\prime}}_{t}(\tilde{\mu}_{i+1},\bm{\pi}^{\psi^{\prime}}(\tilde{\mu}_{i+1}))$
23:         end for
24:         Update $\theta$ with gradient descent $\nabla_{\theta}\frac{1}{B}\sum_{i}(q_{i}-Q^{\theta}_{t}(\tilde{\mu}_{i},\bm{\pi}^{\psi}(\tilde{\mu}_{i}))^{2}$
25:         Update $\psi$ with gradient ascent $\nabla_{\psi}\frac{1}{B}\sum_{i}Q^{\theta}_{t}(\tilde{\mu}_{i},\bm{\pi}^{\psi}(\tilde{\mu}_{i}))$
26:         $\theta^{\prime}\leftarrow\alpha\theta+(1-\alpha)\theta^{\prime}$
27:         $\psi^{\prime}\leftarrow\alpha\psi+(1-\alpha)\psi^{\prime}$
28:     end for
29:end for
30:$\bm{\pi}_{n}^{\psi}\leftarrow\bm{\pi}^{\psi^{\prime}}$
Appendix D Experiments
In this section, we provide further analysis, the motivation behind our modeling decisions, and details for making our experiments easily replicable. We use a private cluster with GPUs to run our experiments. Safe-$\text{M}^{3}$-UCRL and $\text{M}^{3}$-UCRL with known dynamics each used 15 minutes of one Intel Xeon Gold 511 CPU core, 32 GB of RAM and one Nvidia GeForce RTX 3090 GPU. Training Safe-$\text{M}^{3}$-UCRL and $\text{M}^{3}$-UCRL with unknown dynamics to produce results in Figure 3 and Figure 4(b) had 24 hours of access to fifty Intel Xeon Gold 511 CPU cores, 64 GB of RAM and fifty Nvidia GeForce RTX 3090 GPUs during the batch execution necessary for training. The evaluation, i.e., generating the results for, e.g., Figure 5(b) and Figure 7 took around 1 hour of one Intel Xeon Gold 511 CPU core, 64 GB of RAM and one Nvidia GeForce RTX 3090 GPU. The only computationally intensive evaluation task was for Figure 5(a) for more than 1 million agents. We had access to Xeon Gold 511 CPU core, 128-256 GB of RAM, for around 8 hours. The implementation was predominantly done in Python packages PyTorch [42] and NumPy [24].
Our experimental workflow has the following structure:
  1. 1.
    Input data preprocessing
    • Estimating the demand distribution for taxi service $\rho_{0}$
    • Estimating passenger’s trip destinations’ likelihood mapping $\Phi(\cdot)$
  2. 2.
    Modeling assumptions, model parameters, and distributions’ representation
    • State-space, action space, noise, reward, safety constraint
    • Mean-field distribution representation
    • Mean-field transition $U(\cdot)$
  3. 3.
    Executing model-based learning protocol (Algorithm 1)
    • Optimizing Section 4
    • Learning unknown transition dynamics using probabilistic ensemble, i.e., statistical estimators $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$
  4. 4.
    Performance evaluation in the infinite regime
  5. 5.
    Performance evaluation in the finite regime
D.1 Input Data Preprocessing
We consider taxi trajectories collected in Shenzhen’s extended city center with the geographical area spanning from 114.015 to 114.14 degrees longitude and from 22.5 to 22.625 degrees latitude. We have access to the trajectories of five full working weeks (Monday to Sunday) collected between 18th January 2016 and 25th September 2016. We restrict ourselves to evening peak hours between 19:00 and 22:00. We represent probability distributions by discretizing the space into $k\times k$ unit grid with $k=25$ where each cell, $C_{ij}=[\frac{i}{k},\frac{i+1}{k}\rangle\times[\frac{j}{k},\frac{j+1}{k}\rangle$ for $i,j\in\{0,\ldots,k-1\}$, represents a square neighborhood of around $550\times 550$ meters on the city map. We represent taxi service demand distribution $\rho_{0}\in\mathcal{P}(\mathcal{S})$ as a $k\times k$ matrix where entries $[\rho_{0}]_{ij},i,j\in\{0,\ldots,k-1\}$ represent a probability of a trip originating in the neighborhood $C_{ij}$. The probabilities are estimated as an average over the considered period and kept constant during each step $t$ of the learning protocol (see Section D.3). To smooth out the demand distribution and remove possible noise in the raw data, we apply 2-dimensional median smoothing with a window equal to 3 and show the output in Figure 1(a). If the passenger’s trip originates at the state $s\in\mathcal{S}$ the likelihood of its destinations is defined by the mapping $\Phi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ which we use in Section D.2 to define sequential transitions by first executing passenger trips followed by vehicle repositioning (see Figure 1). We flatten the $k\times k$ space grid into $k^{2}$-dimensional vector and represent $\Phi(\cdot)$ as a $k^{2}\times k^{2}$ matrix. The entries $[\Phi]_{ij}$ with $i,j\in\{0,\ldots,k^{2}-1\}$ represent the likelihood of a passenger’s trip that originated in the neighborhood $i$ ending in the neighborhood $j$. The likelihoods are estimated as an average over the considered period and kept constant. Figure 4 shows $\Phi$ column average, i.e., the trip destinations’ likelihood given an arbitrary trip origin.
Figure 4: Passenger’s trip destinations’ likelihood during evening peak hours. The image shows the trip destinations’ likelihood given an arbitrary trip origin (see Section D.1). We observe that the trips to most residential areas are almost equally likely and dispersed across the entire study region, which is consistent with the intuition that some residents commute back home (outside of the city center) after a work day, while others commute towards the city center for, e.g., leisure activities.
D.2 Modeling Assumptions and Model Parameters
We represent our area of interest as a two-dimensional unit square, i.e., the state-space $\mathcal{S}=[0,1]^{2}$, and assume that taxis can move freely using repositioning actions taken from $\mathcal{A}=[-1,1]^{2}$. If the action takes a taxi outside of the state-space borders, we project it back onto the border. Since the fleet is operating in a noisy traffic environment, repositioning usually cannot be executed perfectly. We model the noise $\varepsilon_{n,t}\sim\text{TN}(0,\sigma^{2}I_{2})$ by a Gaussian with a known variance $\sigma^{2}$ truncated at the borders of $\mathcal{S}$ and $I_{2}$ is the $2\times 2$ unit matrix. We use a standard deviation $\sigma=0.0175$ to represent that the passenger will be dropped off with a 68% probability within a circle with a 240-meter radius of the requested destination or with 95% probability within a circle with a 480-meter radius. For simplicity, we assume that every passenger ride lasts fifteen minutes and that the repositioning is executed instantaneously. We assume that the representative agent (or multiple representative agents) reports the trajectories obtained during the interaction with the environment to the global controller at the end of the day. Therefore, we discretize our modeling horizon in fifteen-minute operational intervals, i.e., $T=12$, and each episode $n$ represents one day. The goal of the service provider is maximizing the profit, which correlates with the amount of satisfied demand. Therefore, we implicitly model the profit by the negative of the Kullback-Leibler divergence between the taxis’ distribution $\mu_{n,t}$ and demand for taxi service $\rho_{0}$, i.e., $r(s_{n,t},\mu_{n,t},a_{n,t})=-D_{KL}(\rho_{0}||\mu_{n,t})$. In other words, $r(\cdot)$ measures the similarity between taxis’ and demand distributions; the closer the distributions higher the profit. In practice, greedy profit maximization is often prevented by imposing service accessibility requirements by regulatory bodies. We use entropic safety constraint $h_{C}(\mu_{n,t})=-\int_{\mathcal{S}}\log(\mu_{n,t}(s))\mu_{n,t}(ds)-C$ to enforce the service accessibility across all residential areas. Therefore, the optimization objective in Section 4 trades off between maximizing the profit $r(\cdot)$ and adhering to accessibility requirements imposed by $h_{C}(\cdot)$. We use matrix $\Phi$ introduced Section D.1 to model sequential transitions by first executing passenger trips followed by vehicle repositioning. Formally, the unknown system dynamics are given by $f(s_{n,t},\mu_{n,t},a_{n,t})=\text{clip}(s^{\Phi}_{n,t}+a_{n,t},0,1)+\varepsilon_{n,t}$, where $s^{\Phi}_{n,t}=\text{sample}(\Phi(s_{n,t}))$. The sampling is done by first determining the destination cell $C_{ij}$ given the origin cell determined by the current state $s_{n,t}$ and origin-destination likelihood matrix $\Phi(\cdot)$. Secondly, we determine the destination state $s^{\Phi}_{n,t}$ by uniform sampling from a cell $C_{ij}$, which is a simplified model of the passenger preferences of the final destination.
Mean-field transition dynamics. During the learning/training phase, we assume that the number of agents $m\rightarrow\infty$ and that these agents induce the mean-field distribution $\mu_{n,t}$ for episode $n$ and step $t$. But, one of the major practical challenges is implementing the mean-field transition function $U(\cdot)$ (see Equation 2). The main difficulties are representing the mean-field distribution $\mu_{n,t}$ and computing the integral in Equation 2 over the state space $\mathcal{S}$. We use discretization to represent the mean-field distribution even though other representations, such as a mixture of Gaussians, are possible. Concretely, we represent the mean-field distribution as $\mu_{n,t}=[\mu_{n,t}]_{ij}$ with $i,j\in\{0,\ldots,k-1\}$ by associating the probability $[\mu_{n,t}]_{ij}=\mathbb{P}\big{[}s_{n,t}\in C_{ij}\big{]}$ of the representative agent residing within each cell $C_{ij}$ during episode $n$ at step $t$. The initial mean-field distribution, $\mu_{n,0}$, follows the uniform distribution to maximize the Shannon entropy, ensuring the safety at the beginning of every episode $n$ at $t=0$, i.e., $h_{C}(\mu_{n,0})\geq 0$. In vehicle repositioning, the mean-field transition dynamics $U(\cdot)$ consist of two sequential steps induced by steps in individual system dynamics $f$. Namely, first, the demand shifts the mean-field distribution, followed by the transition induced by the controller. Formally, the mean-field demand transition is computed as $\mu_{n,t}^{\Phi}=(\mu_{n,t}\cdot p)\times\Phi+\mu_{n,t}\cdot(1-p)$, where $\cdot$ denotes elementwise multiplication, $\times$ denotes matrix multiplication and $p=\min(1,\frac{\rho_{0}}{\mu_{n,t}})$ represents a proportion of occupied taxis. The mean-field controller transition requires computing the integral in Equation 2 for which we use the discrete approximation given the points $c_{ij}$ uniformly chosen from cells $C_{ij}$ for $i,j\in\{0,\ldots,k-1\}$:
$[\mu_{n,t+1}]_{ij}=\sum_{k,l}\mathbb{P}[f(c_{kl},\mu_{n,t}^{\Phi},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))+\varepsilon_{n,t}\in C_{ij}][\mu_{n,t}^{\Phi}]_{kl},$
for the episode $n$ and step $t$. We assume that the noise term $\varepsilon_{n,t}$ is independent across episodes and steps as well as along the two dimensions while the truncation parameters are adjusted relative to the state space borders. Thus, we have the following:
$\begin{split}&\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))+\varepsilon_{n,t}\in C_{ij}\right]\\ &=\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}+\varepsilon_{n,t,x}\in\left[\frac{i}{k},\frac{i+1}{k}\right\rangle\right]\\ &\quad\times\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}+\varepsilon_{n,t,y}\in\left[\frac{j}{k},\frac{j+1}{k}\right\rangle\right]\\ &=\left[\phi\left(\frac{i+1}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}\right)-\phi\left(\frac{i}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}\right)\right]\\ &\quad\cdot\left[\phi\left(\frac{j+1}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}\right)-\phi\left(\frac{j}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}\right)\right],\end{split}$
where $\phi(\cdot)$ is the cumulative distribution function of truncated Gaussian $\text{TN}(0,\sigma^{2})$.
Mean-field transition dynamics in the finite regime In Section D.5, we instantiate a finite number of taxis $m<\infty$ to evaluate the policy performance in a realistic setting. We keep track of taxis’ states $s_{t}^{(l)}$ for every taxi $l\in\{1,\ldots,m\}$ at steps $t=0,\ldots,T$. In this setting, we approximate the mean-field transition $U(\cdot)$ with the normalized two-dimensional histogram $[\mu_{t+1}]_{ij}$ with bins defined by the cells $C_{ij}$ for $i,j\in\{0,\ldots,k-1\}$ given taxis’ next states $s_{t+1}^{(l)}=f(s_{t}^{(l)},\mu_{t}^{\Phi},\pi(s_{t}^{(l)},\mu_{t}^{\Phi}))+\varepsilon_{t}$.
D.3 Model-based Learning Protocol
We use the learning protocol introduced in Algorithm 1 to train Safe-$\text{M}^{3}$-UCRL. For hyperparameters of the model, see Table 1.
To optimize the objective in Section 4 in the subroutine in Line 3, we use MF-BPTT (see Section C.3.1). We parametrize the policy via a fully-connected neural network with two hidden layers of 256 neurons and Leaky-ReLU activations. The output layer returns the agents’ actions using Tanh activation. We use Xavier uniform initialization [21] to randomly initialize weights while we set bias terms to zero. We use 20,000 training epochs with the early stopping if the policy does not improve at least 0.5% within 500 epochs. To prevent gradient explosion, we use L2-norm gradient clipping with max-norm set to 1. Note that in our experiments, a single neural network had enough predictive power to represent the whole policy profile $\bm{\pi}=(\pi_{0},\ldots,\pi_{T-1})$, but using $T$ networks, one for each individual policy $\pi_{t}$ is a natural extension. For further details about hyperparameters, see Table 2. Additionally, note that we parametrize policy by a neural network with Lipschitz continuous activations (Leakly-ReLU and Tanh). In the case of the bounded neural network’s weights, policy satisfies the Lipschitz continuity assumption (see Assumption 2). In practice, we bound the weights via L2-regularization.
We estimate the confidence set of dynamics functions $\mathcal{F}_{n}$ in the subroutine in Line 5 using deep probabilistic ensemble (see Section C.1). We use an ensemble of 10 fully-connected neural networks with two hidden layers of 16 neurons and Leaky-ReLU activations. The output layer returns the mean and covariance of the confidence set; the mean uses linear activation, and the variance uses SoftPlus activation. We minimize the negative log-likelihood (NLL) for each neural network under the assumption of heteroscedastic Gaussian noise as described in [30]. We randomly split the data from the replay buffer into a training set (90%) and a validation set (10%). We use 10,000 training epochs with the early stopping if the performance on the validation set does not improve for at least 0.5% within 100 consecutive epochs. For further details about hyperparameters, see Table 3.
D.4 Performance Evaluation in the Infinite Regime
In this section, we extend the experiments presented in Section 5. We assume that the number of taxis $m\rightarrow+\infty$ and show two important results. First, we show a conservative behavior of Safe-$\text{M}^{3}$-UCRL by training model with Shannon entropy safety constraint (see Remark 2) with the safety threshold $C$ set to $p=50\%$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$. We then evaluate its performance against a much higher safety threshold, i.e., against the safety threshold induced by $p=85\%$. Figure 4(a) shows that the model never violates stricter safety constraint regardless of its training in a weaker setting. Second, we show the data efficiency of the learning protocol (Algorithm 1) by training models with access to one, five, and ten representative agents for data collection. Figure 4(b) shows that the model with unknown dynamics converges to the model trained with known dynamics almost 6 times faster when using ten representative agents instead of one representative agent. It is a very useful result that can be utilized in many applications. For example, in most transportation applications using tens or even hundreds of representative agents often does not cause cost-related issues. On the contrary, it might be more cost-effective to have ten representative agents for a month than one representative agent for a year.
(a) Safe-$\text{M}^{3}$-UCRL conservative behavior
(b) Learning protocol data efficiency
Figure 5: Safe-$\text{M}^{3}$-UCRL conservative behavior and data efficiency. We train 10 randomly initialized policy profiles where each setup uses the entropic safety constraint $h_{C}(\cdot)\geq 0$. We set the safety threshold $C$ as a percentage $p$ of the maximum Shannon entropy, i.e., $C=p\log(k^{2})$ with $k=25$. In (a), the policy profiles trained for satisfying $h_{0.5\log(k^{2})}(\cdot)\geq 0$ never violate $h_{0.85\log(k^{2})}(\cdot)\geq 0$, which shows the conservative behavior of our model. In (b), we show the data efficiency of the learning protocol (Algorithm 1) by comparing learning curves observed when training models that satisfy $h_{0.85\log(k^{2})}(\cdot)\geq 0$ when using one, five and ten representative agents (RA) for data collection. We see that the model trained with 1-RA converges to the performance of the model with known dynamics in around 80 episodes, while it takes 25 and 15 episodes for 5-RA and 10-RA models, respectively. Note that we use log-reward to emphasize learning speeds on a comparable scale.
D.5 Performance Evaluation in the Finite Regime
In this section, we assume a finite number of taxis $m<+\infty$ and approximate the mean-field distribution $\mu_{t}(s)=\lim_{m\xrightarrow{}\infty}\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(s_{t}^{(i)}=s)$ with the empirical distribution as explained in Section D.5. The policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to reposition each of $m$ individual taxis in the fleet. In Figure 5(a), we show the relationship between model performance and the number of taxis in the system. As expected, we observe that increasing the number of taxis leads to better performance due to the increased precision of mean-field distribution $\mu_{t}$ approximation at step $t$. By performing 100 randomly initialized runs for each of various fleet sizes, we see that Safe-$\text{M}^{3}$-UCRL learned with unknown dynamics and applied in the finite regime converges to the solution achieved in the infinite regime with known dynamics. Furthermore, the model performs very well already for a fleet of $10^{4}$ taxis, which is in the order of magnitude of the fleet size that operates in Shenzhen. Concretely, in 2016 the fleet had around 17 thousand taxis, with an increasing trend. We also observe that in the finite regime, the difference in performance between Safe-$\text{M}^{3}$-UCRL trained with known and unknown dynamics becomes insignificant. To showcase the practical usefulness of the algorithm, in Figure 5(b), we instantiate 10,000 taxis and display their positions at step $T=12$ after the final repositioning. We observe that the majority of taxis are repositioned to areas of high demand. At the same time, some of them are sent to residential zones in the northwest and northeast to enforce accessibility. It is important to note that some taxis are repositioned to aquatic areas and areas without infrastructure due to two reasons. First, our model guarantees global safety without explicit guarantees for individual/local safety. Notice that undesirable areas might be avoided by safety constraint shaping, e.g., by setting the weight function for these areas to zero as discussed in Section B.2. Second, the model loses some of its accuracy due to the finite regime approximation errors. In Figure 7, we show only a slight decrease in dispersion when Safe-$\text{M}^{3}$-UCRL finite regime approximation is compared to the infinite regime performance.
Objective from Section 4 achieved in the finite regime by approximating mean-field distribution with empirical distribution
Taxis’ locations after repositioning action
in the finite regime
Figure 6: Safe-$\text{M}^{3}$-UCRL performance in the finite regime. We instantiate a finite number of taxis, each following a policy profile $\bm{\pi}_{n}^{*}$ learned in the infinite regime and satisfying the entropic safety constraint $h_{C}(\cdot)\geq 0$ for $C=0.85\log(k^{2})$ with $k=25$. We perform 100 randomly initialized runs for each of the various fleet sizes. In (a), we see that increasing the number of taxis leads to better performance due to the increased precision of mean-field distribution approximation. Further, we see that Safe-$\text{M}^{3}$-UCRL learned with unknown dynamics and applied in the finite regime converges to the solution achieved in the infinite regime with known dynamics. We also see that the curves for Safe-$\text{M}^{3}$-UCRL trained with unknown and known dynamics almost overlap, i.e., the value of knowing dynamics has little to no insignificance in the finite regime. In (b), we showcase the performance for the realistic number of taxis operating in Shenzhen (10,000 to 20,000). We display the positions of a randomly chosen subset of 1,000 taxis (out of 10,000) at step $T=12$ after the final repositioning. We observe that the majority of taxis are repositioned to areas of high demand, while some of them are sent to residential zones in the northwest and northeast to enforce accessibility.
Safe-$\text{M}^{3}$-UCRL with unknown dynamics
in the infinite regime
Safe-$\text{M}^{3}$-UCRL with unknown dynamics
in the finite regime
Figure 7: Mean-field approximation error in the finite regime. We show the difference in the performance introduced by approximating mean-field distribution with the finite number of taxis as described in Section D.2. In (a), we see that the dispersion of Safe-$\text{M}^{3}$-UCRL in the infinite regime is 96% as elaborated in Figure 2. In (b), the policy profile $\bm{\pi}_{n}^{*}$ trained in the infinite regime is used to control a fleet of 10,000 taxis. We observe a slight decrease in the dispersion from 96% to 92% due to the mean-field distribution approximation errors.
Hyperparameter Value Description
$N$ 200 Number of episodes
$T$ 12 Number of steps
$k$ 25 Number of discretization segments per axis
$\sigma$ 0.0175 Standard deviation of the system noise
$L_{h}$ 0.1 Lipschitz constant in Section 4
$\lambda$ 1 Log-barrier hyperparameter in Section 4
# of representative agents 1 to 10 By default 1, but in some experiments, we use more
Table 1: Learning protocol hyperparameters
Hyperparameter Value Description
# of hidden layers 2
# of neurons 256 Number of neurons per hidden layer
hidden activations Leaky-ReLU
output activation Tanh
$\alpha$ $10^{-4}$ Learning rate
$w$ $5\cdot 10^{-4}$ Weight decay
initialization Xavier uniform
bias initialization 0
$n$ 20,000 Number of epochs
early stopping 0.5% If not improved after 500 epochs
Table 2: Policy hyperparameters
Hyperparameter Value Description
# of ensemble members 10
# of hidden layers 2
# of neurons 16 Number of neurons per hidden layer
hidden activations Leaky-ReLU
mean output activation Linear
covariance output activation SoftPlus
$\alpha$ $10^{-4}$ Learning rate
$w$ $5\cdot 10^{-4}$ Weight decay
$\beta$ 1 Assumption 4 hyperparameter
initialization Xavier uniform
bias initialization 0
$n$ 10,000 Number of epochs
early stopping 0.5% If not improved for 100 consecutive epochs
train-validation split 90%-10% We use a validation set for early stopping
replay buffer size 10-100 By default 100
batch size 8-128
Table 3: Probabilistic ensemble hyperparameters